If an air their is number density n=10-19cm-3 and T= 300k . Justify weather this air is plasma or not ?


1
Expert's answer
2020-07-30T10:38:53-0400

Let us calculate how many particles normal air at 300 K contain within 1 cubic centimetre:


"N=nN_A=\frac{m}{\Mu}N_A=\frac{\rho V}{\Mu}N_A,\\\space\\
N=\frac{1.161\cdot1}{29}6.022\cdot10^{23}=2.41\cdot10^{22}\text{ cm}^{-3}."


This is how many particles air would contain in 1 cm3 at 300 K. Therefore, the given substance is plasma.